3.Calculate the volumeoccupied by 5.0 moles of nitrogen gas (N2) at 28°C and 2.05 atm in a disinfectant


aerosol spray.



2. Calculate the number of moles of oxygen(O2) in a 35.0 L ventilator cylinder at STP.


1
Expert's answer
2021-12-08T12:09:02-0500

"1)\\P=2.05atm"

"V="


"n=5moles"

"R=0.0821. \frac{atmL}{molK}"


"T=(28+273)K=301K"


"PV=nRT\implies\>V=\frac{nRT}{P}"


"V=\frac{5×0.0821×301}{2.05}"


"=60.27L"





"2)\\\>P=1atm\\V=35L\\n=\\R=0.0821\>\frac{atmL}{molK}"


"T=273"


"PV=nRT\implies\>n=\frac{PV}{RT}"

"n=\frac{1×35}{0.0821×273}\\\\=1.5616moles"